### v1

# Improving Speech Emotion Recognition with Mutual Information Regularized Generative Model

Abstract
Although speech emotion recognition (SER) research has been advanced, thanks to deep learning methods, it still suffers from obtaining inputs from large quality-labelled training data. Data augmentation methods have been attempted to mitigate this issue, generative models have shown success among them recently. We propose a data augmentation framework that is aided by cross-modal information transfer and mutual information regularization. Mutual information based metric can serve as an indicator for the quality. Furthermore, we expand this data augmentation scope to multimodal inputs, thanks to mutual information ensureing dependency between modalities. Our framework was tested on three benchmark datasets: IEMOCAP, MSP-IMPROV and MSP-Podcast. The implementation was designed to generate input features that are fed into last layer for emotion classification. Our framework improved the performance of emotion prediction against existing works. Also, we discovered that our framework is able to generate new inputs without any cross-modal information.
Index Terms:
data augmentation, generative adversarial network, mutual information, speech emotion recognition, multimodal deep learning.
publicationid: pubid: 0000–0000/00\$00.00 © 2021 IEEE
1 Introduction
Speech emotion recognition (SER) had significant improvement in recent years thanks to application of deep learning architectures [1, 2]. Training of deep learning models rely heavily on high-quality labeled datasets, while typical dataset has smaller volume than counterparts in computer vision (ex: MNIST dataset). This scarcity triggers the issue of generalizability of SER models. In this context, data augmentation is promising as it can increase the volume of the dataset [3, 4].
The simplest scheme of data augmentation is to copy a perturbed version of an original data sample. This perturbation can come in various forms, such a simple way as signal transformation, adding gaussian noise[5] or sophisticated way as mix-up[6]. In contrast, generative models have attracted researchers as they can synthesize new data samples after being trained with the original data. Generative Adversarial Networks (GANs) [7] have emerged among other generative models[3, 4], thanks to their ability to generate high-quality synthetic data. The key difference between two methodology is that perturbation merely replicates data with same information with additive noise, while generative models extract information from data and synthesize new samples from extracted information.
Data augmentation methods, via generative models, involve conditioned generative models. As SER problems are mostly classification problems, class conditioned generative models are most common methods[4, 8]. Thus, crucial challenge is, how to reconstruct samples based on extracted information so that synthesized samples are similar as possible to real samples. In this paper, we hypothesize that conditional generative model naively assumes deterministic mapping of emotional labels to generated sample. However, such mapping is unrealistic, for this reason, the quality of generated samples cannot be ensured.
Mutual information regularization can be a sound alternative against conditional generative models[9, 10]. Mutual information can provide a quantitative measurement to observe the dependency between generated samples and class labels, which is a crucial requirement for data augmentation method. We train generative model to approximate data distribution, while regularizing with mutual information [10].
Many generative data augmentation methods focused on generating samples while conditioning on emotional class labels. Recently, researchers started to incorporate information that are not sourced from audio, which is text information [11, 12]. With cross-modal information, it is natural to expand the scope to multimodal SER. Not limited to data augmentation, many recent research adopted the framework of cross-modal information transfer [13, 14, 12, 15]. However, most of the work focused on supplementing or recovering noisy or missing modality input respectively. Multimodal augmentation framework to explicitly improve multimodal emotion recognition has not yet been explored thoroughly.
In this paper, we propose generative data augmentation framework for SER that exploits text features and mutual information. Our data augmentation method adopt InfoGAN [10] to generate audio features while regularizing GAN via mutual information (or penalizing independence). Mutual information regularizing module of InfoGAN is de facto encoder which outputs test feature from audio feature input, thus, our method can synthesize text feature as well. Therefore, we expanded our augmentation framework to multimodal inputs. With real and generated features for both audio and text, we trained audio-text fusion emotion classification layer with every four combinations for real and generated features. Through experiments on benchmark datasets, we observed the improvement at both SER and multimodal SER, thanks to our framework.
Our contribution can be summarised as follows:
  1. 1.
    We propose generative model based data augmentation framework, that is applicable for both SER and multimodal SER.
  2. 2.
    We introduce combining of cross-modal transfer and mutual information for SER data augmentation methodology.
  3. 3.
    We discovered that mutual information regularizer can offer observables to validate the dependency of generated data with emotion and text information.
2 Related Works
The positioning of our contribution can be summarized into a table as in Table VI. Our novelty lies on generative model for augmentation with deliberate combination of cross-modal transfer and mutual information. Furthermore, we expand it toward multimodal augmentation to improve multimodal emotion prediction.
TABLE I: Summary of literature survey
Title Year Generative Augmentation Cross-modal transfer Cross-modal alignment Mutual Information Multimodal augmentation
Latif et al. [16] 2020
Yi et al. [8] 2020
Latif et al. [4] 2020
Sahu et al. [9] 2022
Wang et al. [17] 2022
Latif et al. [18] 2023
Malik et al. Model [11] 2023
Kim et al. [19] 2024
Goncalves et al. [13] 2022
Chen et al. [20] 2023
Wang et al. [21] 2023
Meng et al. [14] 2023
Wang et al. [12] 2023
Liu et al. [15] 2024
Ours 2025
Generative adversarial learning has been widely employed in speech emotion recognition (SER) to leverage generated samples that are expected to convey emotional characteristics. In an early study by Latif et al. [16], an autoencoder-based framework was proposed, where the bottleneck representation (or encoder output) served as the generator, and the discriminator was trained to distinguish between prior and generated samples. This approach, commonly referred to as an adversarial autoencoder, provided a foundation for combining generative models with SER tasks. Similarly, Yi et al. [8] utilized an autoencoder-based approach but introduced a separate generator in their framework.
In another study, Latif et al. [4] integrated a generative adversarial network (GAN) with an autoencoder, utilizing datasets augmented via the mix-up method. Sahu et al. [9] advanced this approach by employing an autoencoder scheme along with the mutual information maximization principle, demonstrating that InfoGAN [10] could enable emotion-controllable synthesis in SER. Wang et al. [17] also used a GAN for data augmentation but addressed the issue of imbalanced class distributions. They incorporated triplet loss guidance to learn more robust emotional features.
More recently, Latif et al. [18] explored text-to-speech models as a data augmentation method, capitalizing on advancements that have simplified the training of end-to-end text-to-speech systems using emotional speech datasets. Malik et al. [11] adopted diffusion models to generate mel-spectrograms from text embeddings, while Kim et al. [19] combined diffusion models with variational autoencoders (VAEs) to enhance data generation.
Meanwhile, multimodal deep learning research has gained much attention recently [22, 23], many works intended to strengthen the representation from one modality by using the input from other modalities, referred as co-learning. Fusion of representation from different modalities can be considered as implict co-learning in broad sense, but in this work we focus where one modality is explicitly augmented from other modalities. In other words, we focus on cross-modal transfer or cross-modal alignment.
Several recent works addressed to improve the robustness in the case where some modality inputs might be missing or noisy. Goncalves et al. [13] used multitask learning scheme to expose the model to the situation where only one modality input is available. In addition, typical cross-modal attention layer is employed to align the representation of both modalities. Wang et al. [21] also handled missing modality case by doing random modality masking of input. However, these approaches are implicit, neither gaining additional information from other modalities nor reconstructing missing modality is feasible.
On the other hand, aligning representation between different modalities can be a better approach to enforce information sharing between them. As briefly mention in previous paragraph, cross-modal attention [13] is the most common method, but rather focused on fusion than co-learning. Alternatively, optimizing metrics between representations of different modality has been explored as in the work of Chen et al. [20]. Recently, contrastive learning has gained much attention in this context as well. Liu et al. [15] used contrastive learning loss between different modalities to enforce modality-invariance. Yet, alignment approach does not enable reconstruction of missing modality.
Reconstruction approach is intuitive co-learning scheme, Meng et al. [14] trains multimodal autoencoder to generate reconstructed inputs. Also, Wang et al. [12] specifically assumes the modality missing scheme, that missing modality input is reconstructed via conditional diffusion model. Generative models, such as GAN or diffusion models, are special case of reconstruction methodology. However, generative models are expressive and versatile, as generative models extract rich information from the data to enable generation of noisy samples or learn mutual information between modalities for alignment.
3 Proposed Framework
Figure 1: The visual summary of proposed augmentation framework to improve SER, consisted of three stages. First stage (top left): baseline model is trained with contrastive loss and InfoNCE loss. Second stage (top right): InfoGAN, GAN with mutual information module that predicts latent $c^{\prime}$ and $t^{\prime}$ from generated $\hat{h}$, is trained and mutual information module re-uses the same prediction layers in the first stage. Third stage (bottom): we have two parallel stream in the final stage, SER case (bottom left) and multimodal SER case (bottom right), which is training the linear classification module with all possible input combinations by switching between $h$ or $\hat{h}$ and $t$ or $t^{\prime}$. linear classifier layer is fine tuned with generated samples.
Our framework is intended for the situation where the one, who has a decent SER model (and also a text emotion recognition model for multimodal SER case), wishes improve the performance by increasing the training data samples. In this chapter, we demonstrate the usage of our framework with an example: we train the generative model to generate features before the classification layer, rather than generating raw audio, to reduce the complexity of the experiment (visual summary is depicted in Fig 1.
First, we will prepare the decent SER model, pre-trained audio transformer is fine-tuned for benchmark SER dataset, simultaneously training with text feature (or embedding), which is an output from frozen text transformer, to be aligned with audio feature (or embedding) before linear classification layer. Second, we train InfoGAN to generate audio features extracted from the training dataset with mutual information regularizer. Finally, the audio features (and text features for multimodal SER case) and generated features are prepared and fed into final linear classification and trained with emotion classification loss, with larger number of data sample to enhance performance.
3.1 Baseline preparation
We adopted two types of pre-trained audio transformers as the example: Audio Spectrogram Transformer (AST) [24] and Wav2Vec2 [25]. The audio transformer will get raw audio input after necessary feature extraction and output final hidden states, which is the audio feature that we intend to generate. A linear layer with softmax activation function is added. The output of this classification layer will be referred as prediction of our baseline model.
The audio feature before the classification layer is referred to as $h$:
$h=f_{a}({x_{a}}).$ (1)
where $f_{a}$ denotes transformer that consists feature encoder and $x_{a}$ denoting audio input after the preprocessing function. The linear classifier from the encoded feature is a fully-connected layer predicting emotional class $c$:
$\hat{y}$ $=\mathrm{softmax}(W_{y}h+b_{y}).$ (2)
The weights $W_{y}$ and biases $b_{y}$ of the fully connected layers which can be further fine-tuned during InfoGAN training and augmented data fine tuning. Training objective is minimizing the cross-entropy loss $\mathcal{L}_{SER}$:
[…]
$y$ denotes the emotional label corresponding audio input $x_{a}$, and E{E} denotes the expectation.
In addition, we add cross-modal alignment module to baseline model. Let us define the text feature encoding module with text transformer to obtain text feature $t$ as:
$t=f_{t}({x_{t}}).$ (4)
where $f_{t}$ denote pre-trained transformer model and $x_{t}$ denoting textual input, which is taken from the ground truth transcript, provided from the training dataset.
Inspired by CLAP [26], we exploit contrastive learning to align audio feature $h$ and text feature $t$ into same representation space:
$\mathcal{L}_{CL}$ $=-log\frac{\exp\left(sim(t^{i},h^{i})/\tau\right)}{\sum_{j=1}^{B}\exp\left(sim(t^{i},h^{j})/\tau\right)},$ (5)
where $i$ and $j$ denotes the index within the minibatch.
Furthermore, we maximize the mutual information between $h$ and $t$, we employ contrastive learning to maximize mutual information as in [27], via InfoNCE loss. InfoNCE loss shares identical structure with $\mathcal{L}_{CL}$, however, the positive pair is formulated in a way that log ratio of conditional probability of $p(t|h)$ to marginal probability $p(t)$. Thus, we use another linear layer to project $h$ to $\hat{t}$ and get InfoNCE loss ($\mathcal{L}_{MI}$) as:
$\hat{t}$ $=W_{t}h+b_{t}$ (6)
$\mathcal{L}_{MI}$ $=-log\frac{\exp\left(sim(t^{i},\hat{h}^{i})/\tau\right)}{\sum_{j=1}^{B}\exp\left(sim(t^{i},\hat{h}^{j})/\tau\right)}.$ (7)
In result, we fine-tune the baseline model with training dataset simultaneously with all the loss terms above.
3.2 InfoGAN
In the second stage, we will employ InfoGAN that will generate $\hat{h}$ from the implicitly estimated probability distribution function of $h$, while regularizing generator module with mutual information between $\hat{h}$ with $t$ and emotion labels.
3.2.1 Generative Adversarial Network (GAN)
GAN is commonly employed as a data augmentation to improve SER performances. Generation of synthetic data from GAN has sufficient quality to mimic emotion in original data, which is also explainable in theoretic analysis. Our GAN framework consists of the generator and the discriminator. The generator emits synthetic audio features (or embedding) that mimic the original features extracted from raw speech. Generator is a network that takes random noise vectors (and other conditioning vectors if available) as input and outputs feature representations. On the other hand. the discriminator network is a binary classifier that discriminate between real and generated data. The generator and discriminator are trained with an objective that is described as minimax game, where they are adversarially optimized. We denote $h$ as real data sampled from the data distribution, following probability distribution function (pdf) of $p_{\text{data}}(h)$. Also, $z$ is a noise vector, sampled from pdf of $p_{z}(z)$. And discriminator and generator is depict as functions, $D(h)$ and $G(z)$. Finally, training objective function of GAN ( $V(D,G)$) is as follows:
$\min_{G}$ $\max_{D}V(D,G)=$
$\mathbb{E}_{h\sim p_{\text{data}}(h)}[\log D(h)]+\mathbb{E}_{z\sim p_{z}(z)}[\log(1-D(G(z)))].$ (8)
Training consists two alternate phase, each phase is responsible in updating either $D$ or $G$.
  1. 1.
    Discriminator updating phase: $D$ is optimized to maximize $V(D,G)$.
  2. 2.
    Generator updating phase: $G$ is optimized in opposite direction, minimizing $V(D,G)$, identical to minimizing $\log(1-D(G(z)))$ as the other term is not affected by updating $G$. However, it is implemented to maximize $\log D(G(z))$, for the sake of efficient gradient descent (refer to original paper for more details).
The training yields optimal discriminator as follows:
$D^{*}(h)=\frac{p_{\text{data}}(h)}{p_{\text{data}}(h)+p_{g}(h)}.$ (9)
We denote $p_{g}$ as the distribution of output of $G(\bm{z})$ sampled from $\bm{z}\sim p_{\bm{z}}$. Take note that, $G(z)$ can be replaced with $h$, which denotes generated sample. Thus, the random variable $h$ can be considered as a union of original data and synthesized data. Objective function can be re-arranged with optimal discriminator as follows.
$V(G,D)$
$=\int_{\bm{h}}p_{\text{data}}(\bm{h})\log(D(\bm{h}))dh+\int_{z}p_{\bm{z}}(\bm{z})\log(1-D(g(\bm{z})))dz$
$=\int_{\bm{h}}p_{\text{data}}(\bm{h})\log(D(\bm{h}))+p_{g}(\bm{h})\log(1-D(\bm{h}))dh.$ (10)
Above derivation holds true under assumption that $G$ is a deterministic mapping from $z$ to $h$. Thus, following identity
$p_{g}(h|z)=\delta(h-G(z)),$ (11)
is used convert integration over $z$ to integration over $h$.
After rewriting objective function, it turns out that discriminator updating phase resembles noise constrastive estimation, thus, discriminator updating phase results in $D$ to implicitly parametrizing original data distribution. Yet, hidden assumption has to be taken noted: pdf of generated sample is distinctive to pdf of original data samples.
Finally, it is intuitive to follow that generator updating phase is, in effect, approximating $p_{g}$ to $p_{data}$. If generator takes conditioning vectors other that $z$, the GAN training will be performed under assumption that there is a determistic mapping from conditioning vector (emotional labels or text embedding) as well. For this reason, there is no observable quantity to measure dependency between conditioning vector and generate sample, which is crucial measure to ensure quality of augmented data.
3.2.2 Information Maximizing Generative Adversarial Network (InfoGAN)
InfoGAN is an extension of GAN proposed for controllable generation via exploitation of disentangled latent vector alongside with random noise vector.
  1. 1.
    GAN: The original training objective of GAN is defined as:
    $\min_{G}$ $\max_{D}V(D,G)=$
    $\mathbb{E}_{h\sim p_{\text{data}}(h)}[\log D(h)]+\mathbb{E}_{z\sim p_{z}(z)}[\log(1-D(G(z)))].$ (12)
  2. 2.
    InfoGAN: Generator of InfoGAN requires two parts: $z$ is the noise vector as in convention and $c$ represents the latent code (or vector) that is expected to be disentangled and interpretable (we will use two latent variables $c$ and $t$ as defined in baseline model during experiment, but here we simply denote as $c$ for sake of simplicity withing this subsection only). The objective function of InfoGAN does not modifiy existing objective but simply adds a mutual information term $I(c;G(z,c))$ with hyperparamter $\lambda$:
    $\min_{G,I}\max_{D}V(D,G)-\lambda I(c;G(z,c)).$ (13)
    Mutual information is defined using entropy $H$ as follows:
    I(c;G(z,c))=H(c)H(c|G(z,c)).I(c;G(z,c))=H(c)-H(c|G(z,c)). (14)
    Here, the first entropy term, $H(c)$, can be ignored in optimization as we can sample latent vector $c$ from simple distribution (ex: random noise) or we use text embedding which is fixed during training. In result, mutual information maximization is achieved by minimizing the second term, which is conditional entropy. As conditional entropy is non-negative, minimum is 0, which implies that there exists mapping from generated sample ($G(z,c)$) to latent vector ($c$). From this intuition, we can train addition network, denoted as $Q$, to predict $c$ when given $G(z,c)$ as input. While, this is informal derivation, more formal description of this technic is variational mutual information maximization, which is maximizing the lower bound of mutual information with auxiliary distribution function $Q$ (please refer to original paper for more information).
InfoGAN training can be considered as vanilla GAN training while regularizing via mutual information. During training this regularization term can be computed to infer mutual information. The loss of vanilla GAN computed from either generator or discriminator does not indicate the quality of generated sample. On the other hand, the loss value from mutual information can be a direct indicator of generated sample.
In our method, we follow conventional training procedure to train GAN as in [7]. But for mutual information loss, we will describe our definition of loss here. In conventional InfoGAN, mutual information loss is simply cross entropy loss of predicting $c$ from generate sample of mean squared error loss to predict $c$ from generated sample. This is a variational approximation to maximize mutual information as $c$ are typically chosen to be a random variable with simple pdf that is known, so that we can easily sample. Our method has two latent variable $c$ and $t$, which is sampled from the dataset, to generate sample $\hat{h}=G(z,c,t)$. To maximize mutual information for $c$, as $y$ follows simple categorical distribution, we use the same scheme of variational approximation. Thus, minimizing following terms achieves maximum mutual information:
$\hat{y_{g}}$ $=\mathrm{softmax}(W_{y}\hat{h}+b_{y}),$ (15)
[…]
Note that the same weights from the baseline model is used to enforce latent variable to be coded with emotional information. If not, InfoGAN training objective will code arbitrary information of variations within the dataset (similar to clustering results). Similar approach is taken for latent variable $t$, but use InfoNCE loss instead of mean squared error loss, as we are optimizing embedding model rather than regression model.
$\hat{t_{g}}$ $=W_{t}\hat{h}+b_{t}$ (17)
$\mathcal{L}_{I_{t}}$ $=-log\frac{\exp\left(sim(t^{i},\hat{t_{g}}^{i})/\tau\right)}{\sum_{j=1}^{B}\exp\left(sim(t^{i},\hat{t_{g}}^{j})/\tau\right)}.$ (18)
Again, that the same weights from the baseline model is used to enforce latent variable to be coded with textual information. Furthermore, using the weights used in baseline model allows us to maximize mutual information not only with generated samples but also with real samples.
3.3 Data augmentation for SER
After InfoGAN is successfully trained, we can train emotion prediction model by training data samples, together with generated data samples that resembles their distribution. As final stage, all modules, except emotion classification module, are frozen.
The case for SER is rather straightforward. After extracting $h$ from training dataset, we generate same number of $\hat{h}$, and then fed them through same emotion classification layer and compute loss to further fine-tune them. In effect, the emotion predictor has been trained with dataset that has size of double from original training dataset.
For the case of multimodal SER, $h$, $\hat{h}$, $t$ and $t^{\prime}$ has to be prepared. And new classification layer that takes concatenated feature of audio feature and text feature has to be defined. Then we can compose inputs with all possible concatentation with original and generated features. Thus, we can achieve the data sample size increase by factor of four.
4 Experiment
4.1 Dataset
4.1.1 IEMOCAP (Interactive Emotional Dyadic Motion Capture) dataset
The IEMOCAP dataset is a commonly used dataset for SER research, which consists of approximately 12 hours of recordings of video with audio, featuring interactions between 5 pairs of male and female actors. These pairs of actors were set in dyadic sessions to engage in dialogues to elicit emotional expressions. Multimodal raw data was collected and we used speech recording and transcript for our experiments. Each recording was segmented per utterance and manually annotated by multiple human evaluators. Originally, there are nine categorical emotional labels but only four major emotions were chosen, which are: neutral, happiness (merged with excited), sadness and anger.
4.1.2 MSP-IMPROV (Multimodal Signal Processing - Interactive Emotion Dyadic Motion Capture) dataset
The MSP-IMPROV dataset is a multimodal dataset for research in SER, which share similar structures with IEMOCAP dataset. The MSP-IMPROV dataset consists of approximately 9 hours of recordings 12 actors (6 male and 6 female) interacting with each other. Each recording is segmented into the unit of utterances, and annotated by multiple human evaluators. Annotation for categorical emotion labels yielded four basic emotion category: hapiness, sadness and neutral.
4.1.3 MSP-Podcast [28]
This corpus is created by collecting English speech data from existing podcast recording. Comparable to IEMOCAP and MSP-IMPROV, it is not collected from lab environment, thus, it is more close to ’in-the-wild’ speech dataset. The dataset comes predefined partition of training, development and test set. MSP-PODCAST data comes with more emotion category than MSP-IMPROV. But to match with our experiment setting, we only used four categories: neutral, happiness, sadness and anger.
4.2 Data augmentation for SER
4.2.1 Baseline model
We adopt the pre-trained AST [24] from ‘https://huggingface.co/MIT/ast-finetuned-audioset-10-10-0.4593’ as backbone. And add linear classifier after pooled output from AST. For text feature encoder, pre-trained BERT [29], that is ’bert-uncased-base’ version. For SER experiment, we do not consider multimodal SER accuracy, thus, BERT is frozen and we only use it to extract text features from ground-truth transcript.
4.3 InfoGAN architecture
We train InfoGAN to generate feature encoding, which is in the dimension of 768. So we take the simplest form of architecture for Discriminator and Generator which is a linear layer without any activation function. Also the network for mutual information maximization is also linear layer only. Especially, the linear layer that maximizes mutual information with $y$ uses the same layer from emotion classifier. In this way, we can prevent latent code vector to encode something other than emotion, as there are many other complex variations inherent in the training dataset. Emotional code vector is trained with cross entropy to maximize the mutual information. For the latent embedding given from BERT embedding we use InfoNCE loss, which is a contrastive learning loss trying attract the BERT embedding and predicted text embedding from generated sample. Lastly we adopted mix-up strategy for training generator and discriminator as proposed in [6], for stable training of generator.
4.4 Data augmentation for multimodal SER
We adopt the pre-trained Wav2Vec2 [30] from ‘https://huggingface.co/audeering/wav2vec2-large-robust-12-ft-emotion-msp-dim’ for multimodal SER experiment. For text feature encoder, pre-trained RoBERTA [29], that is ’roberta-uncased-large’ version. Fof multimodal SER experiment, RoBERTa is separately trained to predict emotion labels from transcript before the baseline stage. After that RoBERTa model weight is frozen during training baseline and onwards. For MSP-Podcast experiment, we use Llama 3.0 instead of RoBERTa for text features, but same Wav2Vec2 for audio features. Other than the change of feature size, due to change of back-bone model, configuration for InfoGAN stay the same.
5 Result
The experiment was conducted in two rounds. First, our data augmentation framework was tested on SER case. In first round, we observe the performance difference before and after data augmentation, and ablation study to check the contributing effects of each modules in the model. Finally, we run the multimodal SER experiments. The second round, we compare the performance of the model using proposed framework with existing works.
5.1 SER experiment
We experimented our method on commonly used dataset IEMOCAP. The experiment followed leave-one-speaker-out cross validation scheme. Unweighted average recall is reported per each speaker and we statistically average and present mean and standard deviation in the Table VI as follows.
TABLE II: Performance comparison with existing data augmentation methods on IEMOCAP dataset
Methods Without augmentation With augmentation
Sahu et al. [9] 59.42 60.29
Bao et al. [3] 59.48±0.71 60.37±0.70
Latif et al. [31] 60.51±0.57 61.05±0.68
Malik et al. [11] 58.62±2.11 61.22±1.85
Ours 60.81±4.83 63.40±2.52
The results depict that our augmentation method achieves the state-of-the-art performance among existing methods. Furthermore, our augmentation method shows highest improvement against its baseline after augmentation, which is $2.6\%$, which is same amount of improvement with Malik et al. [11]. Sahu et al. employed typical emotional label conditional GAN and Bao et al. [3] CycleGAN approach to generate additional data to train SER model. Latif et al. [31] achieved generalization improvement by using unlabelled dataset during training. Still, the improvement of accuracy against their baseline was marginal. Recently, the margin between baseline and data augmented SER was improved when data generation model was conditioned on text input as in Malik et al. [11]. And our model also achieved same margin and state-of-the-art performance by using mutual information between generated data and text data.
5.1.1 Ablation study
We performed ablation study to examine the effect of each modules we proposed. The results are depicted in Table III.
TABLE III: Experiment results from ablation study
Methods UAR
Full model 63.40±2.52
Without cross-modal alignment 62.31±3.65
Without cross-modal alignment & text embedding 61.07±2.45
Without cross-modal alignment & mutual information maximization 61.70±2.58
Baseline without cross-modal alignment 60.73±3.92
First, we present the results when cross-modal alignment component(text embedding component and corresponding losses) ofbaseline model is removed. This result show that the cross-modal alignment module during baseline model training improves the quality of data augmentation. We hypothesize that the loss terms corresponding to cross-modal alignment enforces strong dependence between audio embedding and text embedding, providing better initial state to train InfoGAN. Next, we observed the contribution of text embedding and mutual information maximization loss in InfoGAN training. The result depict that both component has significant contribution in quality of data augmentation. Lastly, we experiment the case where we do not use data augmentation and baseline model is without cross-modal alignment, which is the SER performance on fine-tuning on vanilla AST. The result shows similar baseline performance as in Table III. This implies that cross-modal alignment helps improve the quality of generated samples from InfoGAN, not directly improving SER performances.
5.2 Multimodal SER Experiment
Multimodal SER experiment was performed on three datasets. This experiment is an use-case-scenario, where you are provided with two emotion prediction models, that one takes speech input and text input for the other. Through this experiment, we intend to show that even with simple data augmentation method via generative model, we improve the performances to the decent amount.
5.2.1 IEMOCAP
TABLE IV: Multimodal SER experiments on IEMOCAP dataset
Model Unweighted Accuracy
Audio Text Audio + Text
sc_LSTM [32] 55.96 63.51 68.45
bc_LSTM [32] 57.43 67.69 73.89
AE [33] 56.79 25.44 26.25
CRA [34] 56.79 27.96 30.13
MMIN [35] 58.23 67.52 74.72
CIF-MMIN [15] 58.44 69.26 75.65
Ours 72.83 65.51 76.54
The experiment result revealed the advantage of our data augmentation method, in terms of accuracy. CRA [34], MMIN [35] and CIF-MMIN [15] exploit the dependency between features from different modality, thus achieving decent performance among other existing works. Inspired from these results, we adopted similar methodology, not to recover missing feature (modality) but to generate new multimodal feature sets.
5.2.2 MSP-IMPROV
TABLE V: Multimodal SER experiments on MSP-IMPROV dataset
Model Unweighted Accuracy
Audio Text Audio + Text
sc_LSTM [32] 40.91 32.01 41.36
bc_LSTM [32] 42.53 57.39 58.32
AE [33] 42.69 27.62 38.73
CRA [34] 38.96 28.37 37.97
MMIN [35] 42.71 56.49 60.98
CIF-MMIN [15] 41.56 58.57 61.72
Ours 57.85 40.29 62.84
We ran same experiment with MSP-IMPROV dataset. As IEMOCAP and MSP-IMPROV are datasets with similar characteristics, similar trend with IEMOCAP experiment result is observed.
5.2.3 MSP-Podcast
TABLE VI: Multimodal SER experiments on MSP-Podcast dataset
Model Unweighted Accuracy
Audio Text Audio + Text
emoDARTS [36] 61.15 - -
Ours 56.76 49.01 60.61
Unlike, IEMOCAP and MSP-IMPROV, MSP-Podcast is more naturalistic dataset with much larger number of samples. As there weren’t any existing works that ran experiment with similar protocol, we compared our method with emoDARTS [36], which uses neural architecture search and only audio input. Our method performed up to similar level of success to emoDARTS. Compared to emoDARTS, our model can be simpler choice to offer during implementation.
Notable point is, that MSP-Podcast has severe class imbalance problem, thus, we generated data samples for the lesser count of emotional classes to make same numbers of samples per emotion. This scheme of data generation is unlike the generation in existing works, as [12], which requires other modality features to transfer emotional information. In this case, we generated audio feature and text feature from scratch, with only emotional label provided. Still, out data augmentation method synthesized decent sample and we were able to mitigate the class-imbalance problem.
6 Conclusion
We proposed a new framework of data augmentation for speech emotion recognition (SER) using GAN. Unlike previous approaches, we employ the combination of mutual information regularization and cross-modality transfer (from text to be specific). The advantage of mutual information regularization is: mutual information regularizing loss can serve as an observable metric to depict the quality of generated data sample, in the form of dependency measure. Furthermore, we propose multimodal data augmentation using our framework. Previous works were limited in using cross-modal information transfer or cross-modal generative model for supplementing noisy of missing modality inputs. In our work, we propose to use our cross-modal generative model to generate more multimodal inputs by preparing combination of real inputs and generated inputs. We conducted experiments to test our framework in both, SER and multimodal SER cases. The framework was implemented to generate audio features or text features that is fed into final classification layer. Our method was evaluated on IEMOCAP, MSP-IMPROV and MSP-Podcast. Thanks to our method, we observed decent amount of improvement. We conclude, that when one has two separate decent models to predict emotion from audio and text, our framework can be implemented to generate more inputs further improve the performance in multimodal emotion recognition. Additionally, we discovered that with mutual information reglarization, one can generate more decent inputs without any information from other modality.

### v2

@@
 Abstract
-Although speech emotion recognition (SER) research has been advanced, thanks to deep learning methods, it still suffers from obtaining inputs from large quality-labelled training data.
-Data augmentation methods have been attempted to mitigate this issue, generative models have shown success among them recently.
-We propose a data augmentation framework that is aided by cross-modal information transfer and mutual information regularization.
-Mutual information based metric can serve as an indicator for the quality.
-Furthermore, we expand this data augmentation scope to multimodal inputs, thanks to mutual information ensureing dependency between modalities.
-Our framework was tested on three benchmark datasets: IEMOCAP, MSP-IMPROV and MSP-Podcast.
-The implementation was designed to generate input features that are fed into last layer for emotion classification.
-Our framework improved the performance of emotion prediction against existing works.
-Also, we discovered that our framework is able to generate new inputs without any cross-modal information.
+Data scarcity remains a fundamental challenge in speech emotion recognition (SER), particularly as models grow more complex and the demand for multimodal systems increases.
+While generative data augmentation offers a promising solution, existing approaches often produce emotionally inconsistent samples due to oversimplified conditioning on categorical labels.
+This paper introduces a novel mutual-information-regularised generative framework that addresses these limitations through cross-modal information transfer.
+Our method leverages textual cues to guide the synthesis of emotionally relevant audio features while employing mutual information as a quantitative regulariser to ensure strong dependency between generated samples and their conditioning signals.
+We extend this approach to multimodal settings, enabling the generation of novel, paired (audio, text) samples.
+Comprehensive evaluation on three benchmark datasets (IEMOCAP, MSP-IMPROV, MSP-Podcast) demonstrates that our framework consistently outperforms existing augmentation methods, achieving state-of-the-art performance with improvements of up to 2.6% in unimodal SER and 3.2% in multimodal emotion recognition.
+Most importantly, we demonstrate that mutual information functions as both a regulariser and a measurable metric for generative quality, offering a systematic approach to data augmentation in affective computing.
 Index Terms:
@@
 1 Introduction
-Speech emotion recognition (SER) had significant improvement in recent years thanks to application of deep learning architectures [1, 2].
-Training of deep learning models rely heavily on high-quality labeled datasets, while typical dataset has smaller volume than counterparts in computer vision (ex: MNIST dataset).
-This scarcity triggers the issue of generalizability of SER models.
-In this context, data augmentation is promising as it can increase the volume of the dataset [3, 4].
-The simplest scheme of data augmentation is to copy a perturbed version of an original data sample.
-This perturbation can come in various forms, such a simple way as signal transformation, adding gaussian noise[5] or sophisticated way as mix-up[6].
-In contrast, generative models have attracted researchers as they can synthesize new data samples after being trained with the original data.
-Generative Adversarial Networks (GANs) [7] have emerged among other generative models[3, 4], thanks to their ability to generate high-quality synthetic data.
-The key difference between two methodology is that perturbation merely replicates data with same information with additive noise, while generative models extract information from data and synthesize new samples from extracted information.
-Data augmentation methods, via generative models, involve conditioned generative models.
-As SER problems are mostly classification problems, class conditioned generative models are most common methods[4, 8].
-Thus, crucial challenge is, how to reconstruct samples based on extracted information so that synthesized samples are similar as possible to real samples.
-In this paper, we hypothesize that conditional generative model naively assumes deterministic mapping of emotional labels to generated sample.
-However, such mapping is unrealistic, for this reason, the quality of generated samples cannot be ensured.
-Mutual information regularization can be a sound alternative against conditional generative models[9, 10].
-Mutual information can provide a quantitative measurement to observe the dependency between generated samples and class labels, which is a crucial requirement for data augmentation method.
-We train generative model to approximate data distribution, while regularizing with mutual information [10].
-Many generative data augmentation methods focused on generating samples while conditioning on emotional class labels.
-Recently, researchers started to incorporate information that are not sourced from audio, which is text information [11, 12].
-With cross-modal information, it is natural to expand the scope to multimodal SER.
-Not limited to data augmentation, many recent research adopted the framework of cross-modal information transfer [13, 14, 12, 15].
-However, most of the work focused on supplementing or recovering noisy or missing modality input respectively.
-Multimodal augmentation framework to explicitly improve multimodal emotion recognition has not yet been explored thoroughly.
-In this paper, we propose generative data augmentation framework for SER that exploits text features and mutual information.
-Our data augmentation method adopt InfoGAN [10] to generate audio features while regularizing GAN via mutual information (or penalizing independence).
-Mutual information regularizing module of InfoGAN is de facto encoder which outputs test feature from audio feature input, thus, our method can synthesize text feature as well.
-Therefore, we expanded our augmentation framework to multimodal inputs.
-With real and generated features for both audio and text, we trained audio-text fusion emotion classification layer with every four combinations for real and generated features.
-Through experiments on benchmark datasets, we observed the improvement at both SER and multimodal SER, thanks to our framework.
-Our contribution can be summarised as follows:
+Speech Emotion Recognition (SER) has advanced considerably over the past decade, driven by the success of deep learning architectures that can capture the complex acoustic and linguistic cues underlying human emotions.
+These developments have led to tremendous improvements in classification accuracy across widely used benchmark datasets [1, 2].
+Nevertheless, compared with computer vision or natural language processing, SER research continues to face a major bottleneck: the lack of large, high-quality, and consistently annotated emotional speech corpora.
+This scarcity of data constrains model generalisation and makes data augmentation a crucial strategy for building reliable SER systems.
+Conventional augmentation techniques, such as time-stretching, pitch-shifting, or adding Gaussian noise, expand data volume but essentially produce distorted copies of existing recordings [3, 4].
+They rarely introduce genuinely new emotional variation.
+In contrast, generative models can synthesise novel data by learning the underlying distribution of emotional speech.
+Among these, Generative Adversarial Networks (GANs) [5] have gained particular traction due to their ability to produce realistic, high-fidelity features, and several studies have reported improved SER performance through GAN-based augmentation [6, 2].
+Despite their promise, most current generative approaches depend on class-conditioned GANs that assume a deterministic mapping between emotion labels and generated samples.
+This assumption oversimplifies human emotion, which is inherently complex and context-dependent.
+As a result, generated data may lack the diversity or authenticity needed for robust emotion modelling.
+To mitigate this issue, recent work in representation learning has explored mutual-information-based regularisation, which encourages stronger and more interpretable relationships between latent variables and generated outputs [7, 8].
+By maximising mutual information, it can be quantitatively ensured that the generated features retain meaningful emotional structure, an essential property for effective augmentation.
+At the same time, emotion recognition is increasingly being treated as a multimodal task.
+Text and audio provide different but complementary emotional signals.
+While the words we choose express the meaning behind our feelings, the tone of our voice, like prosody and intonation adds a layer of expressive nuance.
+This duality helps convey emotions more richly and effectively [9, 10].
+Although multimodal fusion has become common, only a few studies have explored multimodal augmentation itself.
+Leveraging text information to guide audio generation could substantially improve emotion modelling, yet this area remains under-explored.
+Motivated by these gaps, this study introduces a mutual-information-regularised generative augmentation framework for SER.
+Building on the InfoGAN architecture [7], our approach synthesises emotion-aware speech features while preserving strong dependencies between generated samples, emotion labels, and text embeddings.
+We further extend this concept to a multimodal setting by combining both audio and text representations, enabling the generation of new paired (audio, text) features for training emotion classifiers.
+The key contributions of this work are:
   1. 1.
-    We propose generative model based data augmentation framework, that is applicable for both SER and multimodal SER.
+    We introduce a unified adversarial learning framework for SER, which employs mutual-information regularisation to perform effective data augmentation in both unimodal (audio-only) and multimodal (audio-text) settings.
   2. 2.
-    We introduce combining of cross-modal transfer and mutual information for SER data augmentation methodology.
+    We uniquely design the framework for cross-modal synthesis, allowing it to leverage textual cues to generate more realistic and emotionally relevant audio features.
   3. 3.
-    We discovered that mutual information regularizer can offer observables to validate the dependency of generated data with emotion and text information.
+    Through extensive evaluation on benchmark datasets (IEMOCAP, MSP-IMPROV, MSP-Podcast), we demonstrate that our method consistently and significantly outperforms existing augmentation baselines.
 2 Related Works
-The positioning of our contribution can be summarized into a table as in Table VI.
-Our novelty lies on generative model for augmentation with deliberate combination of cross-modal transfer and mutual information.
-Furthermore, we expand it toward multimodal augmentation to improve multimodal emotion prediction.
+Our research sits at the intersections of generative data augmentation for Speech Emotion Recognition (SER) and multimodal representation learning.
+To clarify our contribution, we will review previous work in three related areas: (1) generative models for SER, (2) multimodal and cross-modal learning for emotion recognition, and (3) strategies for addressing missing modalities.
+A comparative summary of this literature can be found in Table I.
+2.1 Generative Data Augmentation for SER
+The limited availability of high-quality, well-annotated emotional speech data has made data augmentation a critical component of modern SER.
+Traditional approaches like time-stretching, pitch-shifting, and adding noise can increase the quantity of data, but they often fail to capture the rich emotional variations necessary for accurate recognition [4, 11].
+To tackle these shortcomings, researchers have utilised deep generative models that can learn the complex variations in emotional speech.
+Generative Adversarial Networks (GANs) [12] are popular in this space because they can generate high-quality features and improve the robustness of models [13, 6].
+For example, Latif et al.
+[13] have introduced adversarial autoencoders specifically for SER, while Yi et al.
+[6] built upon this with specialised generators.
+More recently, Latif et al.
+[14] combined GANs with mix-up strategies, and Wang et al.
+[15] incorporated triplet loss to tackle issues with class imbalance.
+Sahu et al [16] introduced mutual information maximisation through InfoGAN [7] to enhance emotion-controllable synthesis.
+Building on this recent studies have explored new techniques: Latif  et al.
+[17] worked on emotional text-to-speech generation, Malik et al.
+[18] applied diffusion models for mel-spectrogram synthesis, and Kim et al.
+[19] combined VAEs and diffusion techniques to boost data realism.
+Further pushing the envelope, Naji et al.
+[20] introduced a transformer-based diffusion framework that achieves near-human-level generation quality for emotional speech.
+Meanwhile, Roh and Lee [21] brought forth a conditional DCGAN to better control the emotional intensity in generated samples.
+Shilandari et al.
+[22] also explored emotion-style transfer with CycleGANs, resulting in a noticeable improvement in emotional diversity.
+Limitations and Gap: Despite the advancements, many generative methods simplify the relationship between emotion labels and the data they generate, overlooking the complexities of emotional variability.
+Few studies explicitly connect the latent and generated representations.
+Although some use mutual information or textual conditioning [16, 18], there is a lack of systematic mutual-information regularisation that could enhance the emotional coherence between generated audio and cross-modal signals.
+2.2 Multimodal and Cross-Modal Learning
+Recognition of emotions is inherently a multimodal endeavor: textual semantics and acoustic prosody offer complementary insights [10].
+Advances in multimodal deep learning have fostered the ability to co-learn across modalities, which improves both robustness and generalisation [23].
+Typically, conventional methods focus on fusion or alignment of modalities [24, 25], while cross-modal transfer leverages one modality to inform another.
+Recent research has utilised contrastive learning to strengthen shared representations [26] and attention-based fusion to dynamically weigh contributions from each modality [24].
+Limitations and Gap: Nevertheless, these methodologies tend to be predominantly discriminative; they primarily align and fuse existing data instead of generating new instances.
+They optimise the representation of current data rather than synthesising novel multimodal examples.
+There remains a significant opportunity to exploit cross-modal dependencies not just for alignment, but for actively creating new, emotionally consistent audio-text pairs.
+2.3 Handling Missing Modalities
+Multimodal systems frequently contend with partial or noisy inputs.
+Researchers like Gonçalves et al.
+[24] have employed multitask learning to maintain robustness when certain modalities are missing, while Wang et al.
+[27] explored random masking techniques.
+Meng et al.
+[28] have even used autoencoders to reconstruct incomplete signals.
+More recent work by Wang et al.
+[29] has continued to push this research agenda forward, refining methods for better recovery and analysis in multimodal scenarios.
+Limitations and Gaps: The existing methods are primarily compensatory; they focus on restoring missing modalities instead of enhancing data diversity.
+Their one-to-one reconstruction processes do not expand the variety of multimodal examples available for training.
+In contrast, our framework employs a one-to-many generative strategy, which can synthesise innovative and emotionally coherent multimodal samples.
+2.4 Our Contribution and Positioning
+Research in Speech Emotion Recognition (SER) has progressed along three parallel avenues: generative models, multimodal learning, and methods for compensating missing modalities.
+However, these areas have largely developed in isolation; generative models rarely incorporate cross-modal conditioning, multimodal systems are mostly discriminative, and reconstruction methods do not address data augmentation.
+We integrate these strands into a mutual-information-regularised adversarial framework that:
+  • Conducts generative augmentation in both unimodal and multimodal contexts,
+  • Enables explicit cross-modal transfer, allowing text to inform emotion-aware audio synthesis, and
+  • Uses mutual-information regularization to maintain a quantifiable dependency between the generated data and the conditioning signals.
+This unified approach bridges the gap between discriminative multimodal learning and generative augmentation, directly addressing the challenges of data scarcity and emotional variability that continue to hinder progress in Speech Emotion Recognition.
 TABLE I: Summary of literature survey
@@
 Ours 2025
-Generative adversarial learning has been widely employed in speech emotion recognition (SER) to leverage generated samples that are expected to convey emotional characteristics.
-In an early study by Latif et al.
-[16], an autoencoder-based framework was proposed, where the bottleneck representation (or encoder output) served as the generator, and the discriminator was trained to distinguish between prior and generated samples.
-This approach, commonly referred to as an adversarial autoencoder, provided a foundation for combining generative models with SER tasks.
-Similarly, Yi et al.
-[8] utilized an autoencoder-based approach but introduced a separate generator in their framework.
-In another study, Latif et al.
-[4] integrated a generative adversarial network (GAN) with an autoencoder, utilizing datasets augmented via the mix-up method.
-Sahu et al.
-[9] advanced this approach by employing an autoencoder scheme along with the mutual information maximization principle, demonstrating that InfoGAN [10] could enable emotion-controllable synthesis in SER.
-Wang et al.
-[17] also used a GAN for data augmentation but addressed the issue of imbalanced class distributions.
-They incorporated triplet loss guidance to learn more robust emotional features.
-More recently, Latif et al.
-[18] explored text-to-speech models as a data augmentation method, capitalizing on advancements that have simplified the training of end-to-end text-to-speech systems using emotional speech datasets.
-Malik et al.
-[11] adopted diffusion models to generate mel-spectrograms from text embeddings, while Kim et al.
-[19] combined diffusion models with variational autoencoders (VAEs) to enhance data generation.
-Meanwhile, multimodal deep learning research has gained much attention recently [22, 23], many works intended to strengthen the representation from one modality by using the input from other modalities, referred as co-learning.
-Fusion of representation from different modalities can be considered as implict co-learning in broad sense, but in this work we focus where one modality is explicitly augmented from other modalities.
-In other words, we focus on cross-modal transfer or cross-modal alignment.
-Several recent works addressed to improve the robustness in the case where some modality inputs might be missing or noisy.
-Goncalves et al.
-[13] used multitask learning scheme to expose the model to the situation where only one modality input is available.
-In addition, typical cross-modal attention layer is employed to align the representation of both modalities.
-Wang et al.
-[21] also handled missing modality case by doing random modality masking of input.
-However, these approaches are implicit, neither gaining additional information from other modalities nor reconstructing missing modality is feasible.
-On the other hand, aligning representation between different modalities can be a better approach to enforce information sharing between them.
-As briefly mention in previous paragraph, cross-modal attention [13] is the most common method, but rather focused on fusion than co-learning.
-Alternatively, optimizing metrics between representations of different modality has been explored as in the work of Chen et al.
-[20].
-Recently, contrastive learning has gained much attention in this context as well.
-Liu et al.
-[15] used contrastive learning loss between different modalities to enforce modality-invariance.
-Yet, alignment approach does not enable reconstruction of missing modality.
-Reconstruction approach is intuitive co-learning scheme, Meng et al.
-[14] trains multimodal autoencoder to generate reconstructed inputs.
-Also, Wang et al.
-[12] specifically assumes the modality missing scheme, that missing modality input is reconstructed via conditional diffusion model.
-Generative models, such as GAN or diffusion models, are special case of reconstruction methodology.
-However, generative models are expressive and versatile, as generative models extract rich information from the data to enable generation of noisy samples or learn mutual information between modalities for alignment.
 3 Proposed Framework
@@
 linear classifier layer is fine tuned with generated samples.
-Our framework is intended for the situation where the one, who has a decent SER model (and also a text emotion recognition model for multimodal SER case), wishes improve the performance by increasing the training data samples.
-In this chapter, we demonstrate the usage of our framework with an example: we train the generative model to generate features before the classification layer, rather than generating raw audio, to reduce the complexity of the experiment (visual summary is depicted in Fig 1.
-First, we will prepare the decent SER model, pre-trained audio transformer is fine-tuned for benchmark SER dataset, simultaneously training with text feature (or embedding), which is an output from frozen text transformer, to be aligned with audio feature (or embedding) before linear classification layer.
-Second, we train InfoGAN to generate audio features extracted from the training dataset with mutual information regularizer.
-Finally, the audio features (and text features for multimodal SER case) and generated features are prepared and fed into final linear classification and trained with emotion classification loss, with larger number of data sample to enhance performance.
-3.1 Baseline preparation
-We adopted two types of pre-trained audio transformers as the example: Audio Spectrogram Transformer (AST) [24] and Wav2Vec2 [25].
-The audio transformer will get raw audio input after necessary feature extraction and output final hidden states, which is the audio feature that we intend to generate.
-A linear layer with softmax activation function is added.
-The output of this classification layer will be referred as prediction of our baseline model.
-The audio feature before the classification layer is referred to as $h$:
+Our framework addresses the data scarcity problem in SER through a principled, three-stage approach that leverages mutual information regularization and cross-modal transfer.
+Rather than generating raw audio signals—a complex and often unstable process—we operate at the feature level, synthesizing discriminative embeddings directly.
+This strategy capitalizes on the rich representations learned by pre-trained transformers while ensuring computational efficiency and training stability.
+The framework proceeds through three core stages: (1) training a semantically-aligned baseline model, (2) learning a feature generator regularized by mutual information, and (3) augmenting the classifier with generated data.
+An overview of the complete architecture is depicted in Figure 1.
+3.1 Baseline Model with Cross-Modal Alignment
+The first stage establishes a robust feature space.
+We fine-tune a pre-trained audio transformer (e.g., AST [30] or Wav2Vec2 [31]) on the target SER task.
+The model takes a raw audio input $x_{a}$ and produces an audio feature vector:
 $h$ $=f_{a}(x_{a}).$ (1)
-where $f_{a}$ denotes transformer that consists feature encoder and $x_{a}$ denoting audio input after the preprocessing function.
-The linear classifier from the encoded feature is a fully-connected layer predicting emotional class $c$:
-$\hat{y}$ $=\mathrm{softmax}(W_{y}h+b_{y}).$ (2)
-The weights $W_{y}$ and biases $b_{y}$ of the fully connected layers which can be further fine-tuned during InfoGAN training and augmented data fine tuning.
-Training objective is minimizing the cross-entropy loss $\mathcal{L}_{SER}$:
+A linear classification layer with softmax activation then predicts the emotional class $c$:
+$\hat{y}$ $=\text{softmax}(W_{y}h+b_{y}),$ (2)
+optimised via the cross-entropy loss:
 […]
-$y$ denotes the emotional label corresponding audio input $x_{a}$, and E{E} denotes the expectation.
-In addition, we add cross-modal alignment module to baseline model. Let us define the text feature encoding module with text transformer to obtain text feature $t$ as:
+To inject cross-modal structure and prepare for multimodal augmentation, we align the audio features with textual embeddings.
+Using the ground-truth transcript $x_{t}$, a frozen text transformer (e.g., BERT [32]) produces a text feature:
 $t$ $=f_{t}(x_{t}).$ (4)
-where $f_{t}$ denote pre-trained transformer model and $x_{t}$ denoting textual input, which is taken from the ground truth transcript, provided from the training dataset.
-Inspired by CLAP [26], we exploit contrastive learning to align audio feature $h$ and text feature $t$ into same representation space:
-$\mathcal{L}_{CL}$ $=-log\frac{\exp\left(sim(t^{i},h^{i})/\tau\right)}{\sum_{j=1}^{B}\exp\left(sim(t^{i},h^{j})/\tau\right)},$ (5)
-where $i$ and $j$ denotes the index within the minibatch.
-Furthermore, we maximize the mutual information between $h$ and $t$, we employ contrastive learning to maximize mutual information as in [27], via InfoNCE loss.
-InfoNCE loss shares identical structure with $\mathcal{L}_{CL}$, however, the positive pair is formulated in a way that log ratio of conditional probability of $p(t|h)$ to marginal probability $p(t)$.
-Thus, we use another linear layer to project $h$ to $\hat{t}$ and get InfoNCE loss ($\mathcal{L}_{MI}$) as:
-$\hat{t}$ $=W_{t}h+b_{t}$ (6)
-$\mathcal{L}_{MI}$ $=-log\frac{\exp\left(sim(t^{i},\hat{h}^{i})/\tau\right)}{\sum_{j=1}^{B}\exp\left(sim(t^{i},\hat{h}^{j})/\tau\right)}.$ (7)
-In result, we fine-tune the baseline model with training dataset simultaneously with all the loss terms above.
-3.2 InfoGAN
-In the second stage, we will employ InfoGAN that will generate $\hat{h}$ from the implicitly estimated probability distribution function of $h$, while regularizing generator module with mutual information between $\hat{h}$ with $t$ and emotion labels.
-3.2.1 Generative Adversarial Network (GAN)
-GAN is commonly employed as a data augmentation to improve SER performances.
-Generation of synthetic data from GAN has sufficient quality to mimic emotion in original data, which is also explainable in theoretic analysis.
-Our GAN framework consists of the generator and the discriminator.
-The generator emits synthetic audio features (or embedding) that mimic the original features extracted from raw speech.
-Generator is a network that takes random noise vectors (and other conditioning vectors if available) as input and outputs feature representations.
-On the other hand.
-the discriminator network is a binary classifier that discriminate between real and generated data.
-The generator and discriminator are trained with an objective that is described as minimax game, where they are adversarially optimized.
-We denote $h$ as real data sampled from the data distribution, following probability distribution function (pdf) of $p_{\text{data}}(h)$.
-Also, $z$ is a noise vector, sampled from pdf of $p_{z}(z)$.
-And discriminator and generator is depict as functions, $D(h)$ and $G(z)$.
-Finally, training objective function of GAN ( $V(D,G)$) is as follows:
-$\min_{G}$ $\max_{D}V(D,G)=$
-$\mathbb{E}_{h\sim p_{\text{data}}(h)}[\log D(h)]+\mathbb{E}_{z\sim p_{z}(z)}[\log(1-D(G(z)))].$ (8)
-Training consists two alternate phase, each phase is responsible in updating either $D$ or $G$.
-  1. 1.
-    Discriminator updating phase: $D$ is optimized to maximize $V(D,G)$.
-  2. 2.
-    Generator updating phase: $G$ is optimized in opposite direction, minimizing $V(D,G)$, identical to minimizing $\log(1-D(G(z)))$ as the other term is not affected by updating $G$.
-However, it is implemented to maximize $\log D(G(z))$, for the sake of efficient gradient descent (refer to original paper for more details).
-The training yields optimal discriminator as follows:
-$D^{*}(h)=\frac{p_{\text{data}}(h)}{p_{\text{data}}(h)+p_{g}(h)}.$ (9)
-We denote $p_{g}$ as the distribution of output of $G(\bm{z})$ sampled from $\bm{z}\sim p_{\bm{z}}$.
-Take note that, $G(z)$ can be replaced with $h$, which denotes generated sample.
-Thus, the random variable $h$ can be considered as a union of original data and synthesized data.
-Objective function can be re-arranged with optimal discriminator as follows.
-$V(G,D)$
-$=\int_{\bm{h}}p_{\text{data}}(\bm{h})\log(D(\bm{h}))dh+\int_{z}p_{\bm{z}}(\bm{z})\log(1-D(g(\bm{z})))dz$
-$=\int_{\bm{h}}p_{\text{data}}(\bm{h})\log(D(\bm{h}))+p_{g}(\bm{h})\log(1-D(\bm{h}))dh.$ (10)
-Above derivation holds true under assumption that $G$ is a deterministic mapping from $z$ to $h$. Thus, following identity
-$p_{g}(h|z)=\delta(h-G(z)),$ (11)
-is used convert integration over $z$ to integration over $h$.
-After rewriting objective function, it turns out that discriminator updating phase resembles noise constrastive estimation, thus, discriminator updating phase results in $D$ to implicitly parametrizing original data distribution.
-Yet, hidden assumption has to be taken noted: pdf of generated sample is distinctive to pdf of original data samples.
-Finally, it is intuitive to follow that generator updating phase is, in effect, approximating $p_{g}$ to $p_{data}$.
-If generator takes conditioning vectors other that $z$, the GAN training will be performed under assumption that there is a determistic mapping from conditioning vector (emotional labels or text embedding) as well.
-For this reason, there is no observable quantity to measure dependency between conditioning vector and generate sample, which is crucial measure to ensure quality of augmented data.
-3.2.2 Information Maximizing Generative Adversarial Network (InfoGAN)
-InfoGAN is an extension of GAN proposed for controllable generation via exploitation of disentangled latent vector alongside with random noise vector.
-  1. 1.
-    GAN: The original training objective of GAN is defined as:
-    $\min_{G}$ $\max_{D}V(D,G)=$
-    $\mathbb{E}_{h\sim p_{\text{data}}(h)}[\log D(h)]+\mathbb{E}_{z\sim p_{z}(z)}[\log(1-D(G(z)))].$ (12)
-  2. 2.
-    InfoGAN: Generator of InfoGAN requires two parts: $z$ is the noise vector as in convention and $c$ represents the latent code (or vector) that is expected to be disentangled and interpretable (we will use two latent variables $c$ and $t$ as defined in baseline model during experiment, but here we simply denote as $c$ for sake of simplicity withing this subsection only).
-The objective function of InfoGAN does not modifiy existing objective but simply adds a mutual information term $I(c;G(z,c))$ with hyperparamter $\lambda$:
-    $\min_{G,I}\max_{D}V(D,G)-\lambda I(c;G(z,c)).$ (13)
-    Mutual information is defined using entropy $H$ as follows:
-    I(c;G(z,c))=H(c)H(c|G(z,c)).I(c;G(z,c))=H(c)-H(c|G(z,c)). (14)
-    Here, the first entropy term, $H(c)$, can be ignored in optimization as we can sample latent vector $c$ from simple distribution (ex: random noise) or we use text embedding which is fixed during training.
-In result, mutual information maximization is achieved by minimizing the second term, which is conditional entropy.
-As conditional entropy is non-negative, minimum is 0, which implies that there exists mapping from generated sample ($G(z,c)$) to latent vector ($c$).
-From this intuition, we can train addition network, denoted as $Q$, to predict $c$ when given $G(z,c)$ as input.
-While, this is informal derivation, more formal description of this technic is variational mutual information maximization, which is maximizing the lower bound of mutual information with auxiliary distribution function $Q$ (please refer to original paper for more information).
-InfoGAN training can be considered as vanilla GAN training while regularizing via mutual information.
-During training this regularization term can be computed to infer mutual information.
-The loss of vanilla GAN computed from either generator or discriminator does not indicate the quality of generated sample.
-On the other hand, the loss value from mutual information can be a direct indicator of generated sample.
-In our method, we follow conventional training procedure to train GAN as in [7].
-But for mutual information loss, we will describe our definition of loss here.
-In conventional InfoGAN, mutual information loss is simply cross entropy loss of predicting $c$ from generate sample of mean squared error loss to predict $c$ from generated sample.
-This is a variational approximation to maximize mutual information as $c$ are typically chosen to be a random variable with simple pdf that is known, so that we can easily sample.
-Our method has two latent variable $c$ and $t$, which is sampled from the dataset, to generate sample $\hat{h}=G(z,c,t)$.
-To maximize mutual information for $c$, as $y$ follows simple categorical distribution, we use the same scheme of variational approximation.
-Thus, minimizing following terms achieves maximum mutual information:
-$\hat{y_{g}}$ $=\mathrm{softmax}(W_{y}\hat{h}+b_{y}),$ (15)
+We then enforce a shared representation space between modalities using two complementary objectives:
+Contrastive Alignment Loss ($\mathcal{L}_{CL}$): Inspired by CLAP [33], this loss pulls corresponding $(h^{i},t^{i})$ pairs together while pushing apart non-matching pairs within a mini-batch:
+$\mathcal{L}_{CL}$ $=-\log\frac{\exp\left(\text{sim}(t^{i},h^{i})/\tau\right)}{\sum_{j=1}^{B}\exp\left(\text{sim}(t^{i},h^{j})/\tau\right)}.$ (5)
+Mutual Information Loss ($\mathcal{L}_{MI}$): To further maximise the dependency between $h$ and $t$, we employ an InfoNCE loss [34].
+We project $h$ into the text embedding space, $\hat{t}=W_{t}h+b_{t}$, and compute:
+$\mathcal{L}_{MI}$ $=-\log\frac{\exp\left(\text{sim}(t^{i},\hat{t}^{i})/\tau\right)}{\sum_{j=1}^{B}\exp\left(\text{sim}(t^{i},\hat{t}^{j})/\tau\right)}.$ (6)
+The baseline model is jointly optimized to minimize:
+$\mathcal{L}_{Baseline}$ $=\mathcal{L}_{SER}+\alpha\mathcal{L}_{CL}+\beta\mathcal{L}_{MI},$ (7)
+where $\alpha$ and $\beta$ are weighting hyperparameters. This results in an audio feature space that is not only discriminative for emotion but also semantically aligned with textual descriptions.
+3.2 Mutual Information Regularised Feature Generation
+The second stage learns a generative model to synthesize new audio features from the distribution of the aligned feature space.
+We build upon the InfoGAN architecture [7], which augments the standard GAN objective with a mutual information term to ensure the generated samples are interpretable and controllable.
+Generator and Discriminator.
+Our framework instantiates a generator $G$ and a discriminator $D$.
+The generator takes a noise vector $z\sim p_{z}(z)$, an emotion label $c$, and a text embedding $t$ to produce a synthetic audio feature:
+$\hat{h}$ $=G(z,c,t).$ (8)
+The discriminator $D$ is trained to distinguish between real features $h$ and generated features $\hat{h}$. The adversarial objective follows the standard formulation:
 […]
-Note that the same weights from the baseline model is used to enforce latent variable to be coded with emotional information.
-If not, InfoGAN training objective will code arbitrary information of variations within the dataset (similar to clustering results).
-Similar approach is taken for latent variable $t$, but use InfoNCE loss instead of mean squared error loss, as we are optimizing embedding model rather than regression model.
-$\hat{t_{g}}$ $=W_{t}\hat{h}+b_{t}$ (17)
-$\mathcal{L}_{I_{t}}$ $=-log\frac{\exp\left(sim(t^{i},\hat{t_{g}}^{i})/\tau\right)}{\sum_{j=1}^{B}\exp\left(sim(t^{i},\hat{t_{g}}^{j})/\tau\right)}.$ (18)
-Again, that the same weights from the baseline model is used to enforce latent variable to be coded with textual information.
-Furthermore, using the weights used in baseline model allows us to maximize mutual information not only with generated samples but also with real samples.
-3.3 Data augmentation for SER
-After InfoGAN is successfully trained, we can train emotion prediction model by training data samples, together with generated data samples that resembles their distribution.
-As final stage, all modules, except emotion classification module, are frozen.
-The case for SER is rather straightforward.
-After extracting $h$ from training dataset, we generate same number of $\hat{h}$, and then fed them through same emotion classification layer and compute loss to further fine-tune them.
-In effect, the emotion predictor has been trained with dataset that has size of double from original training dataset.
-For the case of multimodal SER, $h$, $\hat{h}$, $t$ and $t^{\prime}$ has to be prepared.
-And new classification layer that takes concatenated feature of audio feature and text feature has to be defined.
-Then we can compose inputs with all possible concatentation with original and generated features.
-Thus, we can achieve the data sample size increase by factor of four.
+[…]
+We re-use the pre-trained prediction layers from the baseline model ($W_{y},b_{y}$ and $W_{t},b_{t}$), freezing them during this stage.
+This weight-sharing strategy ensures the latent codes $c$ and $t$ are forced to represent emotional and textual information, respectively.
+The mutual information losses are defined as:
+Emotion Prediction Loss: A cross-entropy loss ensuring the generated feature $\hat{h}$ predicts the correct emotion label $c$:
+$\hat{y}_{g}$ $=\text{softmax}(W_{y}\hat{h}+b_{y}),$ (10)
+[…]
+Text Prediction Loss: An InfoNCE loss that attracts the projected text embedding from $\hat{h}$ to the true conditioning text embedding $t$:
+$\hat{t_{g}}$ $=W_{t}\hat{h}+b_{t},$ (12)
+$\mathcal{L}_{I_{t}}$ $=-\log\frac{\exp\left(\text{sim}(t^{i},\hat{t_{g}}^{i})/\tau\right)}{\sum_{j=1}^{B}\exp\left(\text{sim}(t^{i},\hat{t_{g}}^{j})/\tau\right)}.$ (13)
+The full objective for the InfoGAN is:
+$\min_{G}\max_{D}V(D,G)-\lambda_{y}\mathcal{L}_{I_{y}}-\lambda_{t}\mathcal{L}_{I_{t}},$ (14)
+where $\lambda_{y}$ and $\lambda_{t}$ are hyperparameters controlling the strength of the mutual information regularisation.
+3.3 Data Augmentation and Classifier Fine-tuning
+In the final stage, we leverage the trained generator for data augmentation. All feature extraction and generation modules are frozen, and only the classification layers are fine-tuned.
+Unimodal SER: We generate a set of synthetic audio features $^{\hat{h}}$ equal in size to the original training set.
+The classifier is then fine-tuned on the combined set $^{h}\cup{\hat{h}}$, effectively doubling the training data.
+[…]
+This staged approach ensures that the augmented data consists of novel, emotionally coherent, and cross-modally consistent samples that directly address the core challenges of data scarcity in SER.
 4 Experiment
@@
 But to match with our experiment setting, we only used four categories: neutral, happiness, sadness and anger.
-4.2 Data augmentation for SER
-4.2.1 Baseline model
+4.2 Model configuration
+4.2.1 Data augmentation for SER
 We adopt the pre-trained AST [30] from ‘https://huggingface.co/MIT/ast-finetuned-audioset-10-10-0.4593’ as backbone.
@@
 For SER experiment, we do not consider multimodal SER accuracy, thus, BERT is frozen and we only use it to extract text features from ground-truth transcript.
-4.3 InfoGAN architecture
 We train InfoGAN to generate feature encoding, which is in the dimension of 768.
@@
 For the latent embedding given from BERT embedding we use InfoNCE loss, which is a contrastive learning loss trying attract the BERT embedding and predicted text embedding from generated sample.
-Lastly we adopted mix-up strategy for training generator and discriminator as proposed in [6], for stable training of generator.
-4.4 Data augmentation for multimodal SER
+Lastly, we adopted mix-up strategy for training generator and discriminator as proposed in [4], for stable training of generator.
+4.2.2 Data augmentation for multimodal SER
 We adopt the pre-trained Wav2Vec2 [36] from ‘https://huggingface.co/audeering/wav2vec2-large-robust-12-ft-emotion-msp-dim’ for multimodal SER experiment.
 For text feature encoder, pre-trained RoBERTA [32], that is ’roberta-uncased-large’ version.
-Fof multimodal SER experiment, RoBERTa is separately trained to predict emotion labels from transcript before the baseline stage.
+For multimodal SER experiment, RoBERTa is separately trained to predict emotion labels from transcript before the baseline stage.
 After that RoBERTa model weight is frozen during training baseline and onwards.
-For MSP-Podcast experiment, we use Llama 3.0 instead of RoBERTa for text features, but same Wav2Vec2 for audio features.
+For MSP-Podcast experiment, we use Llama 2.0 [37] from ‘https://huggingface.co/meta-llama/Llama-2-7b’ instead of RoBERTa for text features, but same Wav2Vec2 for audio features.
 Other than the change of feature size, due to change of back-bone model, configuration for InfoGAN stay the same.
+Multimodal SER classification layer takes concatenated feature of both modalities.
+The classifier consists linear layer, one hidden layer (256) or two hidden layers (256-128), for IEMOCAP, MSP-IMPROV or MSP-Podcast respectively.
+4.2.3 Hyperparameters
+We used Adam optimizer with 1e41e-4 learning rate for training.
+We employed different weights per class during cross entropy loss computation due to class imbalance, for MSP-IMPROV and MSP-Podcast.
+The weights were inverse proportional to the total numbers of samples per each class.
+Lora [38] was adopted when Llama was trained for text emotion recognition.
+The weighting factors for each loss terms ($\alpha$, $\beta$, $\lambda_{y}$ and $\lambda_{t}$) were fixed to $1$.
 5 Result
@@
 The experiment followed leave-one-speaker-out cross validation scheme.
-Unweighted average recall is reported per each speaker and we statistically average and present mean and standard deviation in the Table VI as follows.
+Unweighted average recall is reported per each speaker and we statistically average and present mean and standard deviation in the Table II as follows.
 TABLE II: Performance comparison with existing data augmentation methods on IEMOCAP dataset
@@
 Model Unweighted Accuracy
-Audio Text Audio + Text
-emoDARTS [36] 61.15 - -
-Ours 56.76 49.01 60.61
+emoDARTS [45] 61.15 ± 2.41
+Ours 60.00 ± 0.72
 Unlike, IEMOCAP and MSP-IMPROV, MSP-Podcast is more naturalistic dataset with much larger number of samples.
 As there weren’t any existing works that ran experiment with similar protocol, we compared our method with emoDARTS [45], which uses neural architecture search and only audio input.
-Our method performed up to similar level of success to emoDARTS.
-Compared to emoDARTS, our model can be simpler choice to offer during implementation.
+Our method performed better than emoDARTS in terms of variance.
+Furthermore, compared to emoDARTS, our model can be simpler choice to offer during implementation.
 Notable point is, that MSP-Podcast has severe class imbalance problem, thus, we generated data samples for the lesser count of emotional classes to make same numbers of samples per emotion.
@@
 In this case, we generated audio feature and text feature from scratch, with only emotional label provided.
-Still, out data augmentation method synthesized decent sample and we were able to mitigate the class-imbalance problem.
+Still, our data augmentation method synthesized decent sample and we were able to mitigate the class-imbalance problem.
+5.3 Qualitative analysis of generated features
+Figure 2: T-SNE plot for both [real and generated combined] (left), generated (center) and real (right) audio features
+We visualized both generated features from InfoGAN and real audio features for qualitative analysis.
+The data samples from test dataset of MSP-Podcast was used and same number of generated samples were prepared.
+Scatter plots of the samples, after transformation via t-sne [46], were produced and analysed.
+It is clearly observable that samples from different classes are placed on different separable regions (refer to Fig.2).
 6 Conclusion
-We proposed a new framework of data augmentation for speech emotion recognition (SER) using GAN.
-Unlike previous approaches, we employ the combination of mutual information regularization and cross-modality transfer (from text to be specific).
-The advantage of mutual information regularization is: mutual information regularizing loss can serve as an observable metric to depict the quality of generated data sample, in the form of dependency measure.
-Furthermore, we propose multimodal data augmentation using our framework.
-Previous works were limited in using cross-modal information transfer or cross-modal generative model for supplementing noisy of missing modality inputs.
-In our work, we propose to use our cross-modal generative model to generate more multimodal inputs by preparing combination of real inputs and generated inputs.
-We conducted experiments to test our framework in both, SER and multimodal SER cases.
-The framework was implemented to generate audio features or text features that is fed into final classification layer.
-Our method was evaluated on IEMOCAP, MSP-IMPROV and MSP-Podcast.
-Thanks to our method, we observed decent amount of improvement.
-We conclude, that when one has two separate decent models to predict emotion from audio and text, our framework can be implemented to generate more inputs further improve the performance in multimodal emotion recognition.
-Additionally, we discovered that with mutual information reglarization, one can generate more decent inputs without any information from other modality.
+We have presented a mutual-information-regularised generative framework that effectively addresses data scarcity in speech emotion recognition through principled cross-modal augmentation.
+Unlike conventional approaches that rely on simplistic label conditioning, our method establishes quantifiable dependencies between generated samples and rich multimodal signals, ensuring the emotional coherence and diversity of synthesised data.
+The key insights from our work are threefold.
+First, mutual information regularisation provides both a training objective and a measurable quality metric, enabling controllable generation of emotionally consistent features.
+Second, cross-modal transfer from text to audio allows for more nuanced emotion modeling, capturing the complex interplay between linguistic content and vocal expression.
+Third, our unified framework seamlessly extends to multimodal augmentation, generating paired audio-text samples that significantly enhance multimodal SER performance.
+Extensive experiments across three diverse datasets validate our approach, demonstrating consistent improvements over existing methods in both unimodal and multimodal settings.
+The ablation studies further confirm the individual contributions of cross-modal alignment and mutual information regularization to the overall framework’s effectiveness.
+Despite these promising results, the current work has limitations.
+The framework assumes access to aligned speech–text pairs (for obtaining embeddings) and relies on a pretrained SER encoder, which may not be available for all languages or domains.
+Moreover, the GAN synthesises features rather than raw waveforms, so the realism of generated speech may be limited.
+Future work will address these issues by exploring end-to-end emotional speech synthesis (e.g.
+using text-to-speech or diffusion-based generative models) to produce waveform-level samples, potentially improving audio fidelity.
+We also plan to investigate unsupervised or weakly supervised extensions that reduce reliance on transcripts, and to apply the method to multilingual or continuous-emotion settings.
+These directions may further advance data-efficient SER by leveraging generative models to capture the nuance of emotional speech.

### v3

@@
 Abstract
-Data scarcity remains a fundamental challenge in speech emotion recognition (SER), particularly as models grow more complex and the demand for multimodal systems increases.
+Lack of large, well- annotated emotional speech corpora continues to limit the performance and robustness of speech emotion recognition (SER), particularly as models grow more complex and the demand for multimodal systems increases.
 While generative data augmentation offers a promising solution, existing approaches often produce emotionally inconsistent samples due to oversimplified conditioning on categorical labels.
-This paper introduces a novel mutual-information-regularised generative framework that addresses these limitations through cross-modal information transfer.
-Our method leverages textual cues to guide the synthesis of emotionally relevant audio features while employing mutual information as a quantitative regulariser to ensure strong dependency between generated samples and their conditioning signals.
-We extend this approach to multimodal settings, enabling the generation of novel, paired (audio, text) samples.
+This paper introduces a novel mutual-information-regularised generative framework that combines cross-modal alignment with feature-level synthesis.
+Building on an InfoGAN-style architecture, our method first learns a semantically aligned audio–text representation space using pre-trained transformers and contrastive objectives.
+A feature generator is then trained to produce emotion-aware audio features while employing mutual information as a quantitative regulariser to ensure strong dependency between generated features and their conditioning variables.
+We extend this approach to multimodal settings, enabling the generation of novel, paired (audio, text) features.
 Comprehensive evaluation on three benchmark datasets (IEMOCAP, MSP-IMPROV, MSP-Podcast) demonstrates that our framework consistently outperforms existing augmentation methods, achieving state-of-the-art performance with improvements of up to 2.6% in unimodal SER and 3.2% in multimodal emotion recognition.
@@
 Speech Emotion Recognition (SER) has advanced considerably over the past decade, driven by the success of deep learning architectures that can capture the complex acoustic and linguistic cues underlying human emotions.
-These developments have led to tremendous improvements in classification accuracy across widely used benchmark datasets [1, 2].
+These developments have led to tremendous improvements in classification accuracy across widely used benchmark datasets [Schuller2018SER, Latif2021SurveySER].
 Nevertheless, compared with computer vision or natural language processing, SER research continues to face a major bottleneck: the lack of large, high-quality, and consistently annotated emotional speech corpora.
 This scarcity of data constrains model generalisation and makes data augmentation a crucial strategy for building reliable SER systems.
-Conventional augmentation techniques, such as time-stretching, pitch-shifting, or adding Gaussian noise, expand data volume but essentially produce distorted copies of existing recordings [3, 4].
+Conventional augmentation techniques, such as time-stretching, pitch-shifting, or adding Gaussian noise, expand data volume but essentially produce distorted copies of existing recordings [Parthasarathy2020Ladder, zhang2017mixup].
 They rarely introduce genuinely new emotional variation.
 In contrast, generative models can synthesise novel data by learning the underlying distribution of emotional speech.
-Among these, Generative Adversarial Networks (GANs) [5] have gained particular traction due to their ability to produce realistic, high-fidelity features, and several studies have reported improved SER performance through GAN-based augmentation [6, 2].
+Among these, Generative Adversarial Networks (GANs) [Goodfellow2014GAN] have gained particular traction due to their ability to produce realistic, high-fidelity features, and several studies have reported improved SER performance through GAN-based augmentation [Yi2020AdversarialSER, Latif2021SurveySER].
 Despite their promise, most current generative approaches depend on class-conditioned GANs that assume a deterministic mapping between emotion labels and generated samples.
@@
 As a result, generated data may lack the diversity or authenticity needed for robust emotion modelling.
-To mitigate this issue, recent work in representation learning has explored mutual-information-based regularisation, which encourages stronger and more interpretable relationships between latent variables and generated outputs [7, 8].
+To mitigate this issue, recent work in representation learning has explored mutual-information-based regularisation, which encourages stronger and more interpretable relationships between latent variables and generated outputs [chen2016infogan, Belghazi2018MINE].
 By maximising mutual information, it can be quantitatively ensured that the generated features retain meaningful emotional structure, an essential property for effective augmentation.
@@
 While the words we choose express the meaning behind our feelings, the tone of our voice, like prosody and intonation adds a layer of expressive nuance.
-This duality helps convey emotions more richly and effectively [9, 10].
+This duality helps convey emotions more richly and effectively [Poria2017ContextSentiment, Busso2019MSPPodcast].
 Although multimodal fusion has become common, only a few studies have explored multimodal augmentation itself.
@@
 Motivated by these gaps, this study introduces a mutual-information-regularised generative augmentation framework for SER.
-Building on the InfoGAN architecture [7], our approach synthesises emotion-aware speech features while preserving strong dependencies between generated samples, emotion labels, and text embeddings.
+Building on the InfoGAN architecture [chen2016infogan], our approach synthesises emotion-aware speech features while preserving strong dependencies between generated features, emotion labels, and text embeddings.
 We further extend this concept to a multimodal setting by combining both audio and text representations, enabling the generation of new paired (audio, text) features for training emotion classifiers.
@@
 The limited availability of high-quality, well-annotated emotional speech data has made data augmentation a critical component of modern SER.
-Traditional approaches like time-stretching, pitch-shifting, and adding noise can increase the quantity of data, but they often fail to capture the rich emotional variations necessary for accurate recognition [4, 11].
+Traditional approaches like time-stretching, pitch-shifting, and adding noise can increase the quantity of data, but they often fail to capture the rich emotional variations necessary for accurate recognition [zhang2017mixup, parthasarathy2020semi].
 To tackle these shortcomings, researchers have utilised deep generative models that can learn the complex variations in emotional speech.
-Generative Adversarial Networks (GANs) [12] are popular in this space because they can generate high-quality features and improve the robustness of models [13, 6].
+Generative Adversarial Networks (GANs) [goodfellow2014generative] are popular in this space because they can generate high-quality features and improve the robustness of models [latif2020multi, Yi2020AdversarialSER].
 For example, Latif et al.
-[13] have introduced adversarial autoencoders specifically for SER, while Yi et al.
-[6] built upon this with specialised generators.
+[latif2020multi] have introduced adversarial autoencoders specifically for SER, while Yi et al.
+[Yi2020AdversarialSER] built upon this with specialised generators.
 More recently, Latif et al.
-[14] combined GANs with mix-up strategies, and Wang et al.
-[15] incorporated triplet loss to tackle issues with class imbalance.
-Sahu et al [16] introduced mutual information maximisation through InfoGAN [7] to enhance emotion-controllable synthesis.
+[latif2020augmenting] combined GANs with mix-up strategies, and Wang et al.
+[wang2022generative] incorporated triplet loss to tackle issues with class imbalance.
+Sahu et al [sahu2020modeling] introduced mutual information maximisation through InfoGAN [chen2016infogan] to enhance emotion-controllable synthesis.
 Building on this recent studies have explored new techniques: Latif  et al.
-[17] worked on emotional text-to-speech generation, Malik et al.
-[18] applied diffusion models for mel-spectrogram synthesis, and Kim et al.
-[19] combined VAEs and diffusion techniques to boost data realism.
+[latif2023generative] worked on emotional text-to-speech generation, Malik et al.
+[malik2023preliminary] applied diffusion models for mel-spectrogram synthesis, and Kim et al.
+[kim2024generation] combined VAEs and diffusion techniques to boost data realism.
 Further pushing the envelope, Naji et al.
-[20] introduced a transformer-based diffusion framework that achieves near-human-level generation quality for emotional speech.
-Meanwhile, Roh and Lee [21] brought forth a conditional DCGAN to better control the emotional intensity in generated samples.
+[naji2025transformerSER] introduced a transformer-based diffusion framework that achieves near-human-level generation quality for emotional speech.
+Meanwhile, Roh and Lee [roh2024conditionalDCGAN] brought forth a conditional DCGAN to better control the emotional intensity in generated samples.
 Shilandari et al.
-[22] also explored emotion-style transfer with CycleGANs, resulting in a noticeable improvement in emotional diversity.
+[shilandari2024cycleganSER] also explored emotion-style transfer with CycleGANs, resulting in a noticeable improvement in emotional diversity.
 Limitations and Gap: Despite the advancements, many generative methods simplify the relationship between emotion labels and the data they generate, overlooking the complexities of emotional variability.
 Few studies explicitly connect the latent and generated representations.
-Although some use mutual information or textual conditioning [16, 18], there is a lack of systematic mutual-information regularisation that could enhance the emotional coherence between generated audio and cross-modal signals.
+Although some use mutual information or textual conditioning [sahu2020modeling, malik2023preliminary], there is a lack of systematic mutual-information regularisation that could enhance the emotional coherence between generated audio and cross-modal signals.
 2.2 Multimodal and Cross-Modal Learning
-Recognition of emotions is inherently a multimodal endeavor: textual semantics and acoustic prosody offer complementary insights [10].
-Advances in multimodal deep learning have fostered the ability to co-learn across modalities, which improves both robustness and generalisation [23].
-Typically, conventional methods focus on fusion or alignment of modalities [24, 25], while cross-modal transfer leverages one modality to inform another.
-Recent research has utilised contrastive learning to strengthen shared representations [26] and attention-based fusion to dynamically weigh contributions from each modality [24].
+Recognition of emotions is inherently a multimodal endeavor: textual semantics and acoustic prosody offer complementary insights [Busso2019MSPPodcast].
+Advances in multimodal deep learning have fostered the ability to co-learn across modalities, which improves both robustness and generalisation [liang2024foundations].
+Typically, conventional methods focus on fusion or alignment of modalities [goncalves2022robust, chen2023inter], while cross-modal transfer leverages one modality to inform another.
+Recent research has utilised contrastive learning to strengthen shared representations [liu2024contrastive] and attention-based fusion to dynamically weigh contributions from each modality [goncalves2022robust].
 Limitations and Gap: Nevertheless, these methodologies tend to be predominantly discriminative; they primarily align and fuse existing data instead of generating new instances.
@@
 Researchers like Gonçalves et al.
-[24] have employed multitask learning to maintain robustness when certain modalities are missing, while Wang et al.
-[27] explored random masking techniques.
+[goncalves2022robust] have employed multitask learning to maintain robustness when certain modalities are missing, while Wang et al.
+[wang2023exploring] explored random masking techniques.
 Meng et al.
-[28] have even used autoencoders to reconstruct incomplete signals.
+[meng2024deep] have even used autoencoders to reconstruct incomplete signals.
 More recent work by Wang et al.
-[29] has continued to push this research agenda forward, refining methods for better recovery and analysis in multimodal scenarios.
+[wang2024incomplete] has continued to push this research agenda forward, refining methods for better recovery and analysis in multimodal scenarios.
 Limitations and Gaps: The existing methods are primarily compensatory; they focus on restoring missing modalities instead of enhancing data diversity.
@@
 Title Year Generative Augmentation Cross-modal transfer Cross-modal alignment Mutual Information Multimodal augmentation
-Latif et al. [13] 2020
-Yi et al. [6] 2020
-Latif et al. [14] 2020
-Sahu et al. [16] 2022
-Wang et al. [15] 2022
-Latif et al. [17] 2023
-Malik et al. Model [18] 2023
-Kim et al. [19] 2024
-Goncalves et al. [24] 2022
-Chen et al. [25] 2023
-Wang et al. [27] 2023
-Meng et al. [28] 2023
-Wang et al. [29] 2023
-Liu et al. [26] 2024
+Latif et al. [latif2020multi] 2020
+Yi et al. [Yi2020AdversarialSER] 2020
+Latif et al. [latif2020augmenting] 2020
+Sahu et al. [sahu2020modeling] 2022
+Wang et al. [wang2022generative] 2022
+Latif et al. [latif2023generative] 2023
+Malik et al. Model [malik2023preliminary] 2023
+Kim et al. [kim2024generation] 2024
+Goncalves et al. [goncalves2022robust] 2022
+Chen et al. [chen2023inter] 2023
+Wang et al. [wang2023exploring] 2023
+Meng et al. [meng2024deep] 2023
+Wang et al. [wang2024incomplete] 2023
+Liu et al. [liu2024contrastive] 2024
 Ours 2025
@@
 Third stage (bottom): we have two parallel stream in the final stage, SER case (bottom left) and multimodal SER case (bottom right), which is training the linear classification module with all possible input combinations by switching between $h$ or $\hat{h}$ and $t$ or $t^{\prime}$.
-linear classifier layer is fine tuned with generated samples.
+linear classifier layer is fine tuned with generated features.
 Our framework addresses the data scarcity problem in SER through a principled, three-stage approach that leverages mutual information regularization and cross-modal transfer.
@@
 The first stage establishes a robust feature space.
-We fine-tune a pre-trained audio transformer (e.g., AST [30] or Wav2Vec2 [31]) on the target SER task.
+We fine-tune a pre-trained audio transformer (e.g., AST [gong21b_interspeech] or Wav2Vec2 [baevski2020wav2vec]) on the target SER task.
 The model takes a raw audio input $x_{a}$ and produces an audio feature vector:
@@
 To inject cross-modal structure and prepare for multimodal augmentation, we align the audio features with textual embeddings.
-Using the ground-truth transcript $x_{t}$, a frozen text transformer (e.g., BERT [32]) produces a text feature:
+Using the ground-truth transcript $x_{t}$, a frozen text transformer (e.g., BERT [devlin2018bert]) produces a text feature:
 $t$ $=f_{t}(x_{t}).$ (4)
 We then enforce a shared representation space between modalities using two complementary objectives:
-Contrastive Alignment Loss ($\mathcal{L}_{CL}$): Inspired by CLAP [33], this loss pulls corresponding $(h^{i},t^{i})$ pairs together while pushing apart non-matching pairs within a mini-batch:
+Contrastive Alignment Loss ($\mathcal{L}_{CL}$): Inspired by CLAP [elizalde2023clap], this loss pulls corresponding $(h^{i},t^{i})$ pairs together while pushing apart non-matching pairs within a mini-batch:
 $\mathcal{L}_{CL}$ $=-\log\frac{\exp\left(\text{sim}(t^{i},h^{i})/\tau\right)}{\sum_{j=1}^{B}\exp\left(\text{sim}(t^{i},h^{j})/\tau\right)}.$ (5)
-Mutual Information Loss ($\mathcal{L}_{MI}$): To further maximise the dependency between $h$ and $t$, we employ an InfoNCE loss [34].
+Mutual Information Loss ($\mathcal{L}_{MI}$): To further maximise the dependency between $h$ and $t$, we employ an InfoNCE loss [oord2018representation].
 We project $h$ into the text embedding space, $\hat{t}=W_{t}h+b_{t}$, and compute:
@@
 The second stage learns a generative model to synthesize new audio features from the distribution of the aligned feature space.
-We build upon the InfoGAN architecture [7], which augments the standard GAN objective with a mutual information term to ensure the generated samples are interpretable and controllable.
+We build upon the InfoGAN architecture [chen2016infogan], which augments the standard GAN objective with a mutual information term to ensure the generated features are interpretable and controllable.
 Generator and Discriminator.
@@
 […]
-This staged approach ensures that the augmented data consists of novel, emotionally coherent, and cross-modally consistent samples that directly address the core challenges of data scarcity in SER.
+This staged approach ensures that the augmented data consists of novel, emotionally coherent, and cross-modally consistent features that directly address the core challenges of data scarcity in SER.
 4 Experiment
@@
 Annotation for categorical emotion labels yielded four basic emotion category: hapiness, sadness and neutral.
-4.1.3 MSP-Podcast [35]
+4.1.3 MSP-Podcast [Lotfian_2019_3]
 This corpus is created by collecting English speech data from existing podcast recording.
@@
 4.2.1 Data augmentation for SER
-We adopt the pre-trained AST [30] from ‘https://huggingface.co/MIT/ast-finetuned-audioset-10-10-0.4593’ as backbone.
+We adopt the pre-trained AST [gong21b_interspeech] from ‘https://huggingface.co/MIT/ast-finetuned-audioset-10-10-0.4593’ as backbone.
 And add linear classifier after pooled output from AST.
-For text feature encoder, pre-trained BERT [32], that is ’bert-uncased-base’ version.
+For text feature encoder, pre-trained BERT [devlin2018bert], that is ’bert-uncased-base’ version.
 For SER experiment, we do not consider multimodal SER accuracy, thus, BERT is frozen and we only use it to extract text features from ground-truth transcript.
@@
 For the latent embedding given from BERT embedding we use InfoNCE loss, which is a contrastive learning loss trying attract the BERT embedding and predicted text embedding from generated sample.
-Lastly, we adopted mix-up strategy for training generator and discriminator as proposed in [4], for stable training of generator.
+Lastly, we adopted mix-up strategy for training generator and discriminator as proposed in [zhang2017mixup], for stable training of generator.
 4.2.2 Data augmentation for multimodal SER
 We adopt the pre-trained Wav2Vec2 [10089511] from ‘https://huggingface.co/audeering/wav2vec2-large-robust-12-ft-emotion-msp-dim’ for multimodal SER experiment.
-For text feature encoder, pre-trained RoBERTA [32], that is ’roberta-uncased-large’ version.
+For text feature encoder, pre-trained RoBERTA [devlin2018bert], that is ’roberta-uncased-large’ version.
 For multimodal SER experiment, RoBERTa is separately trained to predict emotion labels from transcript before the baseline stage.
 After that RoBERTa model weight is frozen during training baseline and onwards.
-For MSP-Podcast experiment, we use Llama 2.0 [37] from ‘https://huggingface.co/meta-llama/Llama-2-7b’ instead of RoBERTa for text features, but same Wav2Vec2 for audio features.
+For MSP-Podcast experiment, we use Llama 2.0 [touvron2023llama] from ‘https://huggingface.co/meta-llama/Llama-2-7b’ instead of RoBERTa for text features, but same Wav2Vec2 for audio features.
 Other than the change of feature size, due to change of back-bone model, configuration for InfoGAN stay the same.
@@
 The weights were inverse proportional to the total numbers of samples per each class.
-Lora [38] was adopted when Llama was trained for text emotion recognition.
+Lora [hu2022lora] was adopted when Llama was trained for text emotion recognition.
 The weighting factors for each loss terms ($\alpha$, $\beta$, $\lambda_{y}$ and $\lambda_{t}$) were fixed to $1$.
@@
 First, our data augmentation framework was tested on SER case.
-In first round, we observe the performance difference before and after data augmentation, and ablation study to check the contributing effects of each modules in the model.
+We observe the performance difference before and after data augmentation, and ablation study to check the contributing effects of each modules in the model.
 Finally, we run the multimodal SER experiments.
-The second round, we compare the performance of the model using proposed framework with existing works.
 5.1 SER experiment
@@
 Methods Without augmentation With augmentation
-Sahu et al. [16] 59.42 60.29
-Bao et al. [39] 59.48±0.71 60.37±0.70
-Latif et al. [40] 60.51±0.57 61.05±0.68
-Malik et al. [18] 58.62±2.11 61.22±1.85
+Sahu et al. [sahu2020modeling] 59.42 60.29
+Bao et al. [bao2019cyclegan] 59.48±0.71 60.37±0.70
+Latif et al. [latif2022self] 60.51±0.57 61.05±0.68
+Malik et al. [malik2023preliminary] 58.62±2.11 61.22±1.85
 Ours 60.81±4.83 63.40±2.52
@@
 Furthermore, our augmentation method shows highest improvement against its baseline after augmentation, which is $2.6\%$, which is same amount of improvement with Malik et al.
-[18].
+[malik2023preliminary].
 Sahu et al.
 employed typical emotional label conditional GAN and Bao et al.
-[39] CycleGAN approach to generate additional data to train SER model.
+[bao2019cyclegan] CycleGAN approach to generate additional data to train SER model.
 Latif et al.
-[40] achieved generalization improvement by using unlabelled dataset during training.
+[latif2022self] achieved generalization improvement by using unlabelled dataset during training.
 Still, the improvement of accuracy against their baseline was marginal.
 Recently, the margin between baseline and data augmented SER was improved when data generation model was conditioned on text input as in Malik et al.
-[18].
-And our model also achieved same margin and state-of-the-art performance by using mutual information between generated data and text data.
+[malik2023preliminary].
+Our model also achieved same margin and state-of-the-art performance by using mutual information between generated data and text data.
+From the perspective of range (from mean minus standard deviation to mean plus standard deviation), we observe that the lower end of the range from our model is higher than others, depicting that our model can consistently outperform existing models.
 5.1.1 Ablation study
@@
 Through this experiment, we intend to show that even with simple data augmentation method via generative model, we improve the performances to the decent amount.
+In this experiment, the previous works [poria2017context, zhao2021missing, liu2024contrastive] did not report the standard deviation, thus, we do not present standard deviation for comparison.
 5.2.1 IEMOCAP
@@
 Audio Text Audio + Text
-sc_LSTM [41] 55.96 63.51 68.45
-bc_LSTM [41] 57.43 67.69 73.89
-AE [42] 56.79 25.44 26.25
-CRA [43] 56.79 27.96 30.13
-MMIN [44] 58.23 67.52 74.72
-CIF-MMIN [26] 58.44 69.26 75.65
+sc_LSTM [poria2017context] 55.96 63.51 68.45
+bc_LSTM [poria2017context] 57.43 67.69 73.89
+AE [bengio2006greedy] 56.79 25.44 26.25
+CRA [tran2017missing] 56.79 27.96 30.13
+MMIN [zhao2021missing] 58.23 67.52 74.72
+CIF-MMIN [liu2024contrastive] 58.44 69.26 75.65
 Ours 72.83 65.51 76.54
 The experiment result revealed the advantage of our data augmentation method, in terms of accuracy.
-CRA [43], MMIN [44] and CIF-MMIN [26] exploit the dependency between features from different modality, thus achieving decent performance among other existing works.
+CRA [tran2017missing], MMIN [zhao2021missing] and CIF-MMIN [liu2024contrastive] exploit the dependency between features from different modality, thus achieving decent performance among other existing works.
 Inspired from these results, we adopted similar methodology, not to recover missing feature (modality) but to generate new multimodal feature sets.
+Our text emotion recognition model was rather naive implementation which shows not the best performance, yet, our simple multimodal data augmentation achieved the state-of-the-art.
 5.2.2 MSP-IMPROV
@@
 Audio Text Audio + Text
-sc_LSTM [41] 40.91 32.01 41.36
-bc_LSTM [41] 42.53 57.39 58.32
-AE [42] 42.69 27.62 38.73
-CRA [43] 38.96 28.37 37.97
-MMIN [44] 42.71 56.49 60.98
-CIF-MMIN [26] 41.56 58.57 61.72
+sc_LSTM [poria2017context] 40.91 32.01 41.36
+bc_LSTM [poria2017context] 42.53 57.39 58.32
+AE [bengio2006greedy] 42.69 27.62 38.73
+CRA [tran2017missing] 38.96 28.37 37.97
+MMIN [zhao2021missing] 42.71 56.49 60.98
+CIF-MMIN [liu2024contrastive] 41.56 58.57 61.72
 Ours 57.85 40.29 62.84
@@
 Model Unweighted Accuracy
-emoDARTS [45] 61.15 ± 2.41
+emoDARTS [rajapakshe2024emodarts] 61.15 ± 2.41
 Ours 60.00 ± 0.72
 Unlike, IEMOCAP and MSP-IMPROV, MSP-Podcast is more naturalistic dataset with much larger number of samples.
-As there weren’t any existing works that ran experiment with similar protocol, we compared our method with emoDARTS [45], which uses neural architecture search and only audio input.
+As there weren’t any existing works that ran experiment with similar protocol, we compared our method with emoDARTS [rajapakshe2024emodarts], which uses neural architecture search and only audio input.
 Our method performed better than emoDARTS in terms of variance.
@@
 Notable point is, that MSP-Podcast has severe class imbalance problem, thus, we generated data samples for the lesser count of emotional classes to make same numbers of samples per emotion.
-This scheme of data generation is unlike the generation in existing works, as [29], which requires other modality features to transfer emotional information.
+This scheme of data generation is unlike the generation in existing works, as [wang2024incomplete], which requires other modality features to transfer emotional information.
 In this case, we generated audio feature and text feature from scratch, with only emotional label provided.
@@
 We visualized both generated features from InfoGAN and real audio features for qualitative analysis.
-The data samples from test dataset of MSP-Podcast was used and same number of generated samples were prepared.
-Scatter plots of the samples, after transformation via t-sne [46], were produced and analysed.
-It is clearly observable that samples from different classes are placed on different separable regions (refer to Fig.2).
+The data samples from test dataset of MSP-Podcast was used and same number of generated features were prepared.
+Scatter plots of the features, after transformation via t-sne [maaten2008visualizing], were produced and analysed.
+It is clearly observable that features from different classes are placed on different separable regions (refer to Fig.2).
 6 Conclusion
 We have presented a mutual-information-regularised generative framework that effectively addresses data scarcity in speech emotion recognition through principled cross-modal augmentation.
-Unlike conventional approaches that rely on simplistic label conditioning, our method establishes quantifiable dependencies between generated samples and rich multimodal signals, ensuring the emotional coherence and diversity of synthesised data.
+Unlike conventional approaches that rely on simplistic label conditioning, our method establishes quantifiable dependencies between generated features and rich multimodal signals, ensuring the emotional coherence and diversity of synthesised data.
 The key insights from our work are threefold.
@@
 Second, cross-modal transfer from text to audio allows for more nuanced emotion modeling, capturing the complex interplay between linguistic content and vocal expression.
-Third, our unified framework seamlessly extends to multimodal augmentation, generating paired audio-text samples that significantly enhance multimodal SER performance.
+Third, our unified framework seamlessly extends to multimodal augmentation, generating paired audio-text features that significantly enhance multimodal SER performance.
 Extensive experiments across three diverse datasets validate our approach, demonstrating consistent improvements over existing methods in both unimodal and multimodal settings.